OpenStudy (anonymous):

A tennis ball machine serves a ball vertically into the air from a height of 2 feet, with an initial speed of 120 feet per second. What is the maximum height, in feet, the ball will attain? Round to the nearest whole foot. pleez help I'm lost.. ):

OpenStudy (anonymous):

yeah i bet since no one serves a tennis ball straight up

OpenStudy (anonymous):

you are trying to find the maximum of \[-16t^2+120t+2\] which is at the vertex q

OpenStudy (anonymous):

first coordinate of the vertex of \(y=ax^2+bx+c\) is always \(-\frac{b}{2a}\) which in your case is \[-\frac{120}{2\times (-16)}\] or \[3.75\]

OpenStudy (anonymous):

so 3.75 is my final answer

OpenStudy (anonymous):

?

OpenStudy (anonymous):

second coordinate is what you get when you replace \(t\) by \(3.75\) in \[-16t^2+120t+2\]

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

use a calculator q

OpenStudy (anonymous):

or use this http://www.wolframalpha.com/input/?i=-16%283.75%29^2%2B120*3.75%2B2

OpenStudy (anonymous):

thank u rue the best eve :D

OpenStudy (anonymous):

lol *you are*

OpenStudy (irishboy123):

the third of the common equations used to model Newtons laws of motion will give a direct answer. ie \[v^2 = u^2 + 2ax\] all three should stick in the memory very easily. http://www.bbc.co.uk/bitesize/higher/physics/mech_matt/analyse_motion/revision/2/